OpenStudy (anonymous):

Find the x-intercepts (if any) for the graph of the quadratic function. 6x2 + 12x + 5 = 0 Give your answers in exact form. Show your work.

OpenStudy (astrophysics):

You can use quadratic equation \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] where a = 6, b = 12, and c = 5

OpenStudy (danjs):

solve what you have for x if there are any solutions

OpenStudy (danjs):

yeah that solv's the general form for x..

OpenStudy (danjs):

ax^2+ bx+c=0

OpenStudy (anonymous):

I dont get it

OpenStudy (anonymous):

I SUCK AT MATH

OpenStudy (astrophysics):

Your equation is in the form of \[ax^2+bx+c=0\] so we can use that formula I wrote above called the quadratic formula and just plug in the values for a, b, and c and solve for x, it will give you your x - intercepts

OpenStudy (anonymous):

6x^2+12x+5=0

OpenStudy (anonymous):

@Astrophysics

OpenStudy (astrophysics):

Yes so a = 6, b = 12, and c = 5 use \[x = \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] you will get two values since it's + and -

OpenStudy (astrophysics):

You're replacing the a,b,c in that formula with the numbers 6,12,5

OpenStudy (anonymous):

I give up :(

OpenStudy (astrophysics):

That's because you didn't even try, just plug in the values into the equation nothing to it